OpenStudy (anonymous):

suppose a bead maker wants to seperate 10,000,000 beads into boxes. she will put 100,000 beads into each ox. into which box would she put the 4,358,216th bead

OpenStudy (anonymous):

Just do it the same way as i showed you in your previous question

OpenStudy (anonymous):

4,358,216/100,000= 43,58 so it's over 43 and less than 44 (100,000x43) 4,300,000 (100,000x44) 4,400,000 the 43th box.

OpenStudy (anonymous):

\[\text{Ceiling}\left[\frac{4358216}{100000}\right]=44 \]